OpenStudy (xmoses1):

Input the equation of the given line. The line including (2, 3) and (2, 5).

OpenStudy (xmoses1):

@SolomonZelman

OpenStudy (xmoses1):

I just need a refresh

OpenStudy (xmoses1):

@phi

OpenStudy (luigi0210):

So you're finding an equation, right?

OpenStudy (solomonzelman):

\[\huge\color{forestgreen}{m=\frac{y_1-y_2}{x_1-x_2}}\]\[(2,3)~~~and~~(2,5)\]\[\huge\color{forestgreen}{\frac{3-5}{2-2}~~~~~~~~->~~~~~~\frac{-2}{0}}\]

OpenStudy (solomonzelman):

The slope is undefined as any time when you have a vertical line. How do I know it's vertical, because the x-coordinates are the same.

OpenStudy (solomonzelman):

Equation? you know the slope is undefined and x-coordinates is 2. it would be just x=2 do you need more explanations on this?

OpenStudy (xmoses1):

Nope i think thats all i needed :) thanks!

OpenStudy (solomonzelman):

Anytime!